OpenStudy (anonymous):

Based on the histogram, how many students were included in the survey? What percentage of the students surveyed babysat for at least eight hours last weekend? Show your work.

HanAkoSolo (jamierox4ev3r):

where is the histogram... lol :D

OpenStudy (anonymous):

I believe it's in white so we can't see it.

OpenStudy (uri):

Looooooooooool. +10 @genius12

OpenStudy (anonymous):

Hold on it's not saving to my files to upload...

OpenStudy (anonymous):

I think the histogram is in white in that link as well. @Ryleighblue @uri

OpenStudy (anonymous):

Ugh. it's uploading .-.

HanAkoSolo (jamierox4ev3r):

according to my mac, its a cookie error... :P lol

OpenStudy (anonymous):

I give up .... math hates me...

HanAkoSolo (jamierox4ev3r):

lol im sorry ryleigh :P

OpenStudy (anonymous):

It's all gewd Jamie :/

OpenStudy (anonymous):

@Ryleighblue Take a picture of the histogram usign the "PrtScn" button near the top right of your keyboard. Paste it in paint, save it, and attach to a post...

OpenStudy (anonymous):

It's my last question toooo :(((

OpenStudy (anonymous):

DEWD I've tried that

OpenStudy (anonymous):

this is all you can see

jimthompson5910 (jim_thompson5910):

The four frequencies are: 10, 15, 25, 15 add them up to get 10+15+25+15 = 65 so there are 65 students total 25+15 = 40 students have worked at least 8 hours (ie 8 hours or more), so the percentage is (40/65)*100 = 0.6153846*100 = 61.53846 % Note: this is approximate

OpenStudy (anonymous):

How do you know? dewed I just submitted it saying I didn't know

jimthompson5910 (jim_thompson5910):

I just explained how I know above

OpenStudy (anonymous):

No like the histo gram isn't even there

jimthompson5910 (jim_thompson5910):

the description of the picture gives the numbers you need

jimthompson5910 (jim_thompson5910):

You don't need to see the actual picture. The text is enough.